find the current flowing in the capacitor. j2 e jsr jion mm 1910 ) Esr
Answer #1

Given:- 554 . منزل tee. mm 2 L00 { se: Apply nodal Anelysis i -. + 03 ñ + heo v(5+j4) و د بلال (۷- ۱۵) ذ ا ها ای ها + 0 = ( ژ+5) 207 +v تكنو + (7-19) دہا16 .0 = looV + 8ov + كه هن+ ایان - 1968 .)۱۹ ز1- १२.537 1 34.52 41-१० - १११. 3.13५ ( १२१.52 A)